is 9y^7 a perfect square? if so what is the square root of it?
Answer by stanbon(75887)

is 9y^7 a perfect square? if so what is the square root of it?
----
No it is not: because 7 is not an even exponent.
---
sqrt(9y^7) = sqrt(9y^6)*sqrt(y) = 3y^3*sqrt(y) or 3y^(7/2)
===========
Cheers,
Stan H.
================